I am so rusty in math, I need to know what percentage of 414 is 47? send me some examples until I get it, please.
Thank you very much.

Answer by Lacey020991(49)

% of 414 is 47
Key words to look for is "of" and "is"
Of means to mulitply
Is means equals
.
So say x is our percent
x * 414 = 47 ----divide 47 by 414 to isolate x
x = .1135 move the decimal place to the right twice
x= 11.35% (approx.)
.
.
Ex. 2
what % of 200 is 52?
.
.
x * 200 = 52 ----divide 52 by 200
x= .26 move decimal place to the right twice
x= 26%
.
.
Ex. 3
what % of 345 is 196.65
.
.
.
x * 345 = 196.65 -----divide 196.65 by 345
x= .57 -------move decimal right 2 places
x= 57%
.
.
.
Ex. 4
63% of what is 126
.
.
.
.
.63 * x = 126 -----you have to change % to deci.(move the decimal 2 places left)
x= 200
so 63% of 200 = 126
.
.
If you need anymore just tell me.